Evaluate 𝒂 so that the sum of the eigen values of 𝑨 is 10. [ π‘Ž 4 βˆ’2 1 3 0 βˆ’6 4 π‘Ž ]


1
Expert's answer
2021-06-21T13:47:04-0400

Given matrix A = a 4 -2

1 3 0

-6 4 a

For finding Eigen values,

Consider characteristic equation,

|A-tI| = 0

t3 - t2(3+2a) - t(16-6a) - (3a2-4a-44) = 0

We know sum of the roots is -(-(3+2a))/1

= 3+2a

Given sum of Eigen values is 10

So, 3+2a = 10

2a = 7

a = 3.5 or 7/2